### v1

# City-on-Web: Real-time Neural Rendering of Large-scale Scenes on the Web

Abstract
NeRF has significantly advanced 3D scene reconstruction, capturing intricate details across various environments. Existing methods have successfully leveraged radiance field baking to facilitate real-time rendering of small scenes. However, when applied to large-scale scenes, these techniques encounter significant challenges, struggling to provide a seamless real-time experience due to limited resources in computation, memory, and bandwidth. In this paper, we propose City-on-Web, which represents the whole scene by partitioning it into manageable blocks, each with its own Level-of-Detail, ensuring high fidelity, efficient memory management and fast rendering. Meanwhile, we carefully design the training and inference process such that the final rendering result on web is consistent with training. Thanks to our novel representation and carefully designed training/inference process, we are the first to achieve real-time rendering of large-scale scenes in resource-constrained environments. Extensive experimental results demonstrate that our method facilitates real-time rendering of large-scale scenes on a web platform, achieving 32FPS at 1080P resolution with an RTX 3060 GPU, while simultaneously achieving a quality that closely rivals that of state-of-the-art methods. Project page: https://ustc3dv.github.io/City-on-Web/.
1 Introduction
NeRF has significantly advanced the field of scene reconstruction, showing an unparalleled ability to capture complex details across diverse environments. Existing works have demonstrated its ability to render small scenes with exceptional quality and performance in real-time [3, 46, 47, 36, 21, 29, 6, 40, 11, 43, 48, 24]. NeRF has also been successfully applied to the rendering of large scenes in offline settings, achieving exceptional visual fidelity and generating intricately detailed results [34, 14, 43, 45, 37].
Despite these successes, real-time neural rendering of large scenes is profoundly challenging due to inherent computational power, memory, and bandwidth limitations across various devices. The challenges mainly include the following aspects. Firstly, traditional NeRF and its variants are resource-intensive, requiring substantial computational power that exceeds what is typically available in such constrained environments. Secondly, the video memory capacity on client devices is frequently limited, imposing significant restrictions on the capability to process and render substantial assets in real-time simultaneously. The substantial resource becomes a critical issue in the real-time rendering of large scenes, necessitating the quick loading and processing of extensive data sets. Lastly, the dependency on data retrieval from remote servers introduces latency, particularly under network bandwidth limitations, further complicating the real-time rendering process. These hurdles collectively form a significant barrier to delivering an uninterrupted and instantaneous visual experience for large-scale scenes.
To address these challenges in real-time rendering of large-scale scenes, we propose our proposed method, City-on-Web. Drawing inspiration from traditional graphics techniques for rendering large-scale scenes [10, 38, 7, 9, 15, 8, 23], we partition the scene into manageable blocks and represent the scene with varying Levels-of-Detail (LOD). We utilize radiance field baking techniques [16, 30], which precompute and store rendering primitives into 3D atlas textures organized in a sparse grid within each block for real-time rendering. However, due to the unavoidable texture resource limitations of shaders, we cannot load all the atlas textures into a single shader. Hence, we represent the scene as a hierarchy of segmented blocks, each rendered by a dedicated shader during rendering.
[…]
2 Related Work
Figure 1: Pipeline of training. (a) We divide the entire scene into different blocks without overlap according to the ground plane. (b) For block $k$ that the ray $\boldsymbol{r}(t)$ passes through, the corresponding shader renders the color $\boldsymbol{c}^{k}$ and opacity $\alpha^{k}$ of each block. We depth sort the blocks that ray traversed, and then render the final result through alpha blending that maintains 3D consistency.
Large-scale Scene Reconstruction.
For radiance field reconstruction of large-scale scenes, a key issue lies in enhancing the model’s representational capacity to adequately capture and render extensive scenes. Block-NeRF [34] and Mega-NeRF [37] address this by adopting a divide-and-conquer strategy, segmenting expansive scenes into smaller blocks, and applying localized NeRF processing to each. This approach significantly improves both the reconstruction quality and the model’s scalability to larger scenes. Switch-NeRF [50] employs a gating network to dispatch 3D points to different NeRF sub-networks. Grid-NeRF [45] utilizes a compact multiresolution feature plane and combines the strengths of smoothness from vanilla NeRF with the local detail capturing ability of feature grid-based methods [26, 5, 31], efficiently reconstructing large scenes with fine details. NeRF++ [48] enhances the reconstruction of unbounded scenes through its innovative multi-spherical representation. On the other hand, Mip-NeRF 360 [1] introduces a scene contraction function to effectively represent scenes that extend to infinity, addressing the challenge of vast spatial extents. F2-NeRF [41] takes this a step further by implementing a warping function for local spaces, ensuring a balance of computational resources and training data across different parts of the scene.
Real-time Rendering.
Early works mainly focus on the real-time rendering of a simple single object. NSVF [21] improves NeRF by introducing a more efficient sparse voxel field, significantly accelerating rendering speed while maintaining high-quality output. KiloNeRF [29] utilizes thousands of small MLPs, each responsible for a tiny scene region, significantly reducing network evaluation time. In contrast, SNeRG [16] leverages pre-computed sparse grids, allowing for direct retrieval of radiance field information without needing network evaluation. Termi-NeRF [28] terminates ray marching in less impactful scene regions, slashing computation time. DONeRF [27] focuses on one sample using a depth oracle network, speeding up rendering while preserving scene quality. Recently, there have been developments that enable real-time rendering of neural radiance fields in small scenes. MERF [30] improves upon SNeRG by utilizing a voxel and triplane hybrid representation to reduce memory usage. MobileNeRF [6] introduces the polygon rasterization rendering pipeline, running NeRF-based novel view synthesis in real-time on mobile devices. BakedSDF [46] bakes volumetric representation into meshes and utilizes spherical harmonics for representing view-dependent color, while NeRF2Mesh [36] iteratively refine both the geometry and appearance of the mesh.
Level of Detail.
Substantial works are devoted to integrating LOD methods into the fabric of traditional computer graphics [7, 8, 9, 15, 20, 23, 17, 22], aiming to streamline rendering processes, reduce memory footprint, bolster interactive responsiveness. Recently, some works begin to apply LOD to the neural implicit reconstruction. NGLoD [32] represents LOD through a sparse voxel octree, where each level of the octree corresponds to a different LOD, allowing for a finer discretization of the surface and more detailed reconstruction as the tree depth increases. Takikawa et al. [33] efficiently encode 3D signals into a compact, hierarchical representation using vector-quantized auto decoder method. BungeeNeRF [43] employs a hierarchical network structure, where the base network focuses on learning a coarse representation of the scene, and subsequent residual blocks are tasked with progressively refining this representation. TrimipRF [18] and LoD-Neus [51] leverage multi-scale triplane and voxel representations to capture scene details at different scales, effectively implementing anti-aliasing to enhance the rendering and reconstruction quality.
3 Background and Motivation
Our exploration begins with an in-depth analysis of two influential works, SNeRG [16] and MERF [30], which have both set benchmarks for real-time rendering of radiance field. SNeRG precomputes and stores a Neural Radiance Fields model in a sparse 3D voxel grid. Each active voxel in SNeRG contains several attributes: density, diffuse color, and specular feature vector that captures view-dependent effects. Additionally, an indirection grid is used to enhance rendering by either indicating empty macroblocks or pointing to detailed texels in a 3D texture atlas. This representation allows real-time rendering on standard laptop GPUs.
The indirection grid assists in raymarching through the sparse 3D grid by passing empty regions and selectively accessing non-zero densities $\sigma_{i}$, diffuse colors $\boldsymbol{c}_{i}$, and feature vectors $\boldsymbol{f}_{i}$ during rendering. Integrating along each ray $\boldsymbol{r}(t)=\boldsymbol{o}+t\boldsymbol{d}$, we compute the sum of the weights, which can be considered as the pixel’s opacity:
${{\alpha}}(\boldsymbol{r})=\sum_{i}w_{i},\ w_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})\alpha_{i},\alpha_{i}=1-e^{-\sigma_{i}\delta_{i}}.$ (1)
The color $\boldsymbol{C}_{d}(\boldsymbol{r})$ and specular feature $\boldsymbol{F}_{s}(\boldsymbol{r})$ along the ray are accumulated using the same weights to compute the final diffuse color and specular feature of ray:
${\boldsymbol{C}}_{d}(\boldsymbol{r})=\sum_{i}w_{i}\boldsymbol{c}_{i},\quad{\boldsymbol{F}}_{s}(\boldsymbol{r})=\sum_{i}w_{i}\boldsymbol{f}_{i}.$ (2)
The step size during ray marching $\delta_{i}$ is equal to the voxel width for an occupied voxel. Subsequently, the accumulated diffuse color and specular feature vector, along with the positional encoding $PE(\cdot)$ of the ray’s view direction, are concatenated to pass through a lightweight deferred MLP $\Phi$ to produce a view-dependent residual color:
${\boldsymbol{C}}(\boldsymbol{r})=\boldsymbol{C}_{d}+\Phi(\boldsymbol{C}_{d},\boldsymbol{F}_{s},PE(\boldsymbol{d})).$ (3)
While SNeRG achieves impressive real-time rendering results, its voxel representation demands substantial memory, which poses limitations for further applications. MERF presents a significant reduction in memory requirements in comparison to extant radiance field methods like SNeRG. By leveraging hybrid low-resolution sparse grid and 2D high-resolution triplanes, MERF optimizes the balance between performance and memory efficiency. Moreover, it incorporates two pivotal strategies to bridge the gap between training and rendering performance. First, MERF simulates finite grid approach during training, querying MLPs at virtual grid corners and applying interpolation to mimic the rendering process closely. Second, MERF simulates quantization during training and employs the straight-through estimator [2], allowing for the simulation of the quantization process while maintaining differentiability, enabling the model to learn and optimize with quantized values without introducing non-differentiable steps during the backward pass, ensuring a smooth training process.
These innovative methods for scene reconstruction offer promising results, but their direct applicability to large scenes remains challenges. MERF’s hybrid voxel-triplane representation, despite being memory-efficient, cannot capture large scenes with intricate details due to its fixed resolution constraint. In our efforts to reconstruct large scenes with high fidelity, dividing them into smaller blocks is a practical solution. This method helps in making sure the reconstruction is detailed and accurate.
However, this approach of dividing scenes means we will end up with more assets to deal during rendering. Web browsers have limits on how much memory they can use, which can make it hard to show these many pieces of large and detailed models. Another challenge is the data transmission during web rendering. It often requires pulling data from servers, and this can be slow due to network delays. As a result, users might experience long wait times when trying to load all the detailed pieces of a large scene at once for rendering. Drawing inspiration from traditional mesh dynamic resource loading and LOD [39], we generate LOD from our segmented reconstruction results which helps in minimizing the load of distant resources. Additionally, by employing a dynamic loading strategy for blocks, we significantly reduce VRAM usage and decrease the wait time for resource transmission.
4 Method
In this section, we present a method for representing and rendering large scenes on the web. Our approach uses hierarchical spatial partitioning and LOD to manage large-scale scenes dynamically (Sec. 4.1). We align the training and rendering stages to ensure consistency (Sec. 4.2), employing multiple shaders and alpha blending for seamless integration of scene blocks. Additionally, the framework includes optimization strategies (Sec. 4.3) and a process for generating LODs (Sec. 4.4) and baking the model (Sec. 4.5) for real-time rendering.
4.1 Large-scale Radiance Field
In the realm of scene reconstruction and rendering, NeRF has made significant strides, achieving compelling results. However, it faces inherent challenges when tasked with representing large scenes on the web. Using a single model to represent such vast scenes proves challenging due to its limited expressiveness, particularly in achieving a detailed and accurate reconstruction. Representing scenes with multiple models in a single resolution increases the overhead during rendering, leading to the loading of numerous resources, which is not conducive to efficient rendering.
To efficiently represent large scenes captured using the fly-through method, we employ hierarchical spatial partitioning combined with LOD to represent the scene. Specifically, we uniformly partition the area into varying blocks within our region of interest on the $xy$ plane (i.e., the ground plane). Each set of partitioned blocks corresponds to a unique LOD level, allowing for dynamic and efficient representation. Within each block, we use a low resolution voxel with a high resolution triplane that stores density, diffuse color, and specular color, to represent the radiance field for web rendering. Additionally, for blocks along the periphery, we utilize the scene contraction function from MERF [30] to account for the data on the boundaries. For internal blocks, we simply adopt settings from the bounded scene.
During the training stage, our scene representation aligns with web rendering. We consistently utilize spatial partitioning to structure the scene into distinct blocks without overlap. However, we train the scene only with the finest level of LOD. Within block $k$ , the following trainable components are introduced: (1) $f^{k}$: an attribute query function which adopts a hash encoding and a mlp decoder that outputs attributes of points such as densities, diffuse color and specular feature (2) $\Phi^{k}$: a deferred MLP accounts for view-dependent effects. (3) $\psi^{k}$: a proposal MLP for sampling.
4.2 Consistent Training and Rendering
It is essential to ensure consistency between the training and rendering stages to achieve high-fidelity rendering results on the web as obtained during training. Due to the limited number of texture units within the web rendering environment, we are compelled to create multiple shaders to render distinct blocks. Specifically, one shader is allocated for storing the texture of an individual block. Each block subsequently renders an image respective to the current camera view. However, a simplistic averaging of these resultant rendering outputs can lead to discernible seams and does not ensure 3D consistency at the inter-block boundaries.
To address this problem, we simulate the process of multiple shaders rendering images and then linearly weighting them together using the volume rendering weights of blocks. For ray $\boldsymbol{r}(t)$, we uniformly sample between the near and far boundaries based on the scene’s bounding box. Then, according to the sample coordinates, we query the corresponding block’s proposal MLP to obtain the sample density, which is transformed into probability distributions along the rays. These probabilities guide a resampling strategy, ensuring a concentration on near-surface features with a few samples.
Figure 2: Visualization of alpha blending during consistent training. The left image shows the rendering results of four separate blocks by their shaders and the combined image after alpha blending, demonstrating how alpha blending correctly handles occlusion relationships. The right image visualizes the weights of rendering sample points before and after alpha blending. This illustration shows the correct occlusion management achieved through depth sorting of blocks followed by alpha blending.
Assuming that the proposal MLP yields samples passing through $\boldsymbol{M}$ blocks with a total of $\boldsymbol{N}$ samples, where each block $k$ has $\boldsymbol{n}_{k}$ samples, we simulate web rendering by performing volume rendering within each block to obtain its individual rendering diffuse color $\boldsymbol{C}_{d}^{k}$, specular feature $\boldsymbol{F}^{k}$ and opacity ${\alpha}^{k}$ according to  Eqs. 1 and 2. Then we get block $k$ final rendering color $\boldsymbol{C}^{k}$ according to Eq. 3. Consequently, for the sake of 3D consistency in rendering, we depth-sort the blocks and apply volume rendering across multiple blocks in sequence, using opacity to generate the volume rendering weights:
$\boldsymbol{C}(\boldsymbol{r})=\sum_{k}^{M}\prod^{k-1}_{j=1}(1-\alpha^{j})\boldsymbol{C}^{k}.$ (4)
Under the Lambertian surface setting where the specular color is zero, the diffuse color and feature vector obtained from volume rendering on the total of $\boldsymbol{N}$ ray samples from Eq. 2 are equal to the results produced by our approach of conducting volume rendering within each block followed by inter-block volume rendering Eq. 4. The proof is given in the supplementary. Thus, our rendering approach maintains 3D consistency and simulates multiple shader rendering on the web as shown in Fig. 2.
During rendering, for the sample point $p_{i}$, we need to access the voxel grid corner and triplane grid corner where the point is located and use interpolation to obtain the sample point’s attributes. During training, we also simulate the voxel and triplane grid points to maintain consistency with the rendering process. By using grid corners’ positions to query the attribute query function $f_{k}$, we obtain the attributes of the grid corners. Through interpolation, we acquire the attributes of the sample points in a manner similar to the rendering pipeline. This simulaition strategy ensures that the values used for volume rendering within each block during training are as closely matched as possible to the values queried from the baked textures.
4.3 Optimization
We use Charbonnier loss [4] for reconstruction and S3IM loss [44] to assist multiple blocks’ model in capturing high-frequency details. Additionally, we use the interlevel loss to provide supervision signal for proposal MLP and distortion loss to reduce floaters like Mip-nerf 360 [1].
Moreover, we random uniform sample points set $\mathcal{P}$ within the bounding box of the scene and apply $L_{1}$ regularization on the alpha to encourage the blocks’ model to predict sparse occupied space:
${\mathcal{L}_{\textrm{sparse}}}=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}|\alpha_{i}|=\frac{1}{|\mathcal{P}|}\sum_{p_{i}\in\mathcal{P}}|1-\sigma_{i}v|,$ (5)
where $v$ is the step size used in real-time rendering. Additionally, we introduce a regularization term for the opacity of the block. This regularization encourages the opacity of the block to be as close to 0 or 1 as possible, implying either full transparency or full opaqueness:
${\mathcal{L}_{\textrm{opacity}}}=-\sum_{k}(\alpha^{k}log(\alpha^{k})+(1-\alpha^{k})log(1-\alpha^{k})).$ (6)
In summary, the overall loss function is:
$\begin{split}{\mathcal{L}_{\textrm{train}}}=\;&{\mathcal{L}_{\textrm{charbonnier}}}+\lambda_{1}{\mathcal{L}_{\textrm{S3IM}}}+\lambda_{2}{\mathcal{L}_{\textrm{interlevel}}}+\\ &\lambda_{3}{\mathcal{L}_{\textrm{distortion}}}+\lambda_{4}{\mathcal{L}_{\textrm{sparse}}}+\lambda_{5}{\mathcal{L}_{\textrm{opacity}}}.\end{split}$ (7)
Method BlockA BlockE Campus Rubble Building
Metric
[…]
NeRFacto
25.28
0.469
0.691
24.61
0.441
0.687
23.47
0.255
0.689
19.02
0.538
0.512
17.70
0.442
0.502
Instant-NGP
24.16
0.595
0.643
22.93
0.599
0.619
24.93
0.380
0.703
20.37
0.629
0.478
17.92
0.625
0.424
Mega-NeRF
25.24
0.566
0.668
25.61
0.468
0.680
22.28
0.472
0.565
23.68
0.558
0.525 20.30
0.526
0.506
Grid-NeRF 25.37
0.478
0.705
24.43
0.483
0.693
Ours 26.07 0.357 0.739 25.07 0.357 0.756 24.73 0.192 0.736 21.32 0.482 0.539 19.71 0.439 0.520
Table 1: Quantitative comparison on five large scene datasets. We report PSNR, LPIPS, and SSIM on the test views. The best and second best results are highlighted. The ’-’ symbol indicates that Grid-NeRF [45] was not evaluated on these datasets due to difficulties in adjusting its training configs beyond the provided configurations for Matrix City [19], resulting in poor performance on other datasets.
4.4 LOD Generation
To guarantee superior rendering quality from elevated perspectives and simultaneously diminish the resource demand for distant scene elements, our method involves generating multiple LOD for the scene. We successfully attained the scene’s lowest LOD in the training stage, ensuring maximal visual fidelity. For the downsampling and integration baking of multiple block models into a unified model, we initially freeze the training of hash encoding and decoder MLP components within these models. Subsequently, we proceed to retrain a new tiny deferred MLP. Following the deferred MLP’s successful retraining, we simulate lower resolution virtual voxels and triplane grid corners within the scenes of these multiple blocks. Lastly, this retrained deferred MLP is refined in collaboration with the network responsible for generating grid corner attributes, thereby optimizing the entire rendering process.
4.5 Baking
After the training stage, we conduct block-based evaluation and store the MLP’s outputs onto discrete grids, which generate segmented rendering resources. This approach facilitates efficient resource management for real-time rendering, as each block’s resources are handled independently. Initially, we render all training rays to collect ray samples. Samples with alpha and weight values above a certain threshold are retained, and samples below the threshold are discarded. The preserved samples are used to mark the adjacent eight grid points as occupied in the binary grids. After generating binary grids to identify occupied voxels, we follow the MERF by baking high-resolution 2D planes and a low-resolution 3D voxel grid in each block. Only the non-empty 3D voxels are stored using a block-sparse format. We downsample the occupancy grid with max-pooling for efficient rendering and skip empty space. To further save storage, we compress textures into the PNG format.
5 Experiments
Figure 3: Qualitative comparisons with existing SOTA methods. By testing different methods across diverse scales and environments, it clearly reveals that our approach excels in recovering finer details and achieves a higher quality of reconstruction.
Figure 4: Comparison to Mega-NeRF in Rubble dataset. The Rubble dataset presents significant variations in lighting. Although our method recovers more detail than Mega-NeRF, our deferred shading model has limited ability to represent view-dependent colors and cannot accurately represent lighting and exposure changes in the data as view-dependent effects. This limitation results in slightly lower PSNR values.
5.1 Experiments setup
Dataset and Metric. Our experiments span across various scales and environments. We have incorporated a real-world urban scene dataset (Campus) and public datasets consisting of real-world rural rubble scenes (Rublle, Building[37] and synthetic city-scale data (BlockA and BlockE in MatrixCity[19]. Our datasets were recorded under uniform, cloudy lighting conditions to minimize variation. To obtain precise pose information, we employed an annular capturing approach, which has a higher overlap rate compared to grid-based capturing methods. Tab. 2 presents an overview of our dataset. To assess the quality and fidelity of our reconstructions, we employ various evaluation metrics, including PSNR, SSIM and LPIPS [49].
Area$(m^{2})$ Images Overlap rate Resolution Altitude
1200$\times$800 6515 $95\%$ 8192$\times$5460 180m
Table 2: Overview of Campus dataset.
Implementations and Baselines. Our method takes posed multi-view images captured using a fly-through camera as input. The training code is built on the nerfstudio framework [35] with tiny-cuda-nn [25] extension. And our real-time viewer is a JavaScript web application whose rendering is implemented through GLS. We set the $512^{3}$ resolution for the voxel and $2048^{2}$ resolution for the triplane within each block. We use a 4-layer MLP with 64 hidden dimensions as an encoder after multi-resolution hash encoding to output density, color, and specular feature. Moreover, a 3-layer MLP with 16 hidden dimensions tiny deferred MLP is developed to predict residual view-dependent color. We sample $16384$ rays per batch and use Adam optimizer with an initial learning rate of $1\times 10^{-2}$ decaying exponentially to $1\times 10^{-3}$. Our model is trained with $50k$ iterations on one NVIDIA A100 GPU. We split the scene into 24 non-overlapping blocks for Campus scene and split other scenes into four blocks. Moreover, we benchmark current real-time rendering methods using three critical parameters: Payload (PL), GPU Memory (VRAM), and Frames Per Scond (FPS). Payload refers to the essential data transmitted during the rendering process. We perform qualitative comparisons between our method and existing SOTA methods for large-scale reconstruction. The Campus dataset is partitioned into six sections based on the reconstruction content. NeRFacto, Instant-NGP, and Grid-NeRF were applied to one of these sections, while in other datasets , they are applied to the entire scene. NeRFacto and Instant-NGP are utilized with the highest hash encoding resolution of $8192^{3}$. Similarly, Mega-NeRF divides the Campus dataset into 24 blocks and other datasets into four blocks. Our experiments focus on a single campus section for comparative analysis with existing real-time rendering methods.
5.2 Results Analysis
We systematically evaluate the performance of both baseline models and our method through qualitative and quantitative comparisons in Tab. 1 and Fig. 3. Notably, our method demonstrates a remarkable enhancement in visual fidelity as reflected by the SSIM and LPIPS metrics, which indicate the extent of detail restoration. Despite a reduction in PSNR compared to the SOTA methods, this is attributable to the fact that LPIPS and SSIM are more sensitive to the recovery of fine details, whereas PSNR mainly measures pixel-wise color accuracy. Our approach achieves higher fidelity reconstructions, revealing finer details due to our partitioned reconstruction strategy. In our method, each ray is evaluated by the deferred MLP only once, as opposed to other methods that evaluate the MLP at every sample point. Consequently, while our method recovers more intricate geometrical detail, it frequently results in color discrepancies with the ground truth image due to unstable lighting conditions and variable exposure, as shown in Fig. 4.
[…]
MobileNeRF
19.99
0.516
0.712
544.8
515.3
223
BakedSDF
22.24
0.627
0.413
712.3
242.1
68
Ours(Block)
24.82 0.741 0.190 128.1 40.6
51
Ours(Total)
526.6
114.4
46
Table 3: Comparison with existing real-time methods. We segment the scene into four blocks, while other methods reconstruct the whole scene. ‘Ours(Block)’ denotes the resource usage for one of these blocks, providing a fair comparison with other methods that process the scene as a single block. ‘Ours (Total)’ represents the cumulative resource usage for the entire scene.
In our evaluation, detailed in Tab. 3, we compare our method with current real-time rendering methods, using one segment of the Campus dataset for testing. These tests, are performed on an NVIDIA RTX 3060 Laptop GPU at a $1920\times 1080$ resolution. The results demonstrate that our method excels in reconstruction quality. We represent each scene block using voxels and triplanes, and store the baked grid attributes as images. This strategy significantly reduces the payload. This reduction notably accelerates resource transmission for web-based rendering applications. However, it is observed that our frame rate during rendering is lower compared to other methods. This is attributed to their rendering pipeline based on mesh rasterization, in contrast to our method, which utilizes volume rendering.
5.3 LOD Result
[…]
LOD3
23.72
0.660
0.306
132.1
40.2
LOD2
24.23
0.682
0.297
841.6
201.7
LOD1
24.73
0.736
0.192
3970.2
1259.6
Table 4: The LOD results on the whole Campus dataset.
Figure 5: Visualization of our LOD result.
Tab. 4 presents the quantitative rendering results at various LOD, along with the corresponding payload and VRAM usage. With increasing LOD, the resources required for rendering significantly decrease. Notably, our method’s lowest LOD level still maintains high fidelity rendering results, as demonstrated in Fig. 5. Our LOD strategy significantly streamlines the management of resource loading on web platforms, which is particularly advantageous in rendering distant blocks, as it requires less VRAM. It is worth noting that the VRAM usage presented in Tab. 4 represents the cumulative memory consumption of all blocks. Our dynamic loading strategy adaptively selects resources to load based on the camera’s field of view and the distance to each block, effectively keeping the peak VRAM usage around 1100MB.
5.4 Ablation Study
In Tab. 5, we conduct an ablation study of our method on one section of Campus dataset. Our model is trained for four blocks with low resolution (voxel resolution of $512^{3}$ and triplane resolution of $2048^{2}$). We also train a single model for the entire scene with high resolution (voxel resolution of $1024^{3}$ and triplane resolution of $4096^{2}$). Due to our non-overlapping scene partitioning strategy, these two representations have the same resolution across the entire scene. However, our reconstruction quality is higher, achieving better FPS and lower GPU memory usage, as well as reduced payload. We also removed our consistent training, including virtual grids and alpha blending during training. As shown in the table, our consistent training significantly improves reconstruction quality.
[…]
high res. model
24.05
0.710
0.201
540.9
147
32
w/o consistent
24.21
0.702
0.281
514.4
110
49
ours
24.82
0.741
0.190
526.6
114
46
Table 5: Ablation Study. The result is test on the one section of Campus dataset.
Figure 6: Limitations of Web-on-City.
6 Conclusion and Discussion
In this work, we introduced City-on-Web, which to our knowledge is the first system that enables real-time neural rendering of large-scale scenes over web using laptop GPUs. Our integration of block partitioning with LOD has significantly reduced the payload on the web platform and improved resource management efficiency. We ensured high-fidelity rendering quality by maintaining consistency between training and rendering. Extensive experiments have also fully proved the effectiveness of City-on-Web.
Limitation & Future Work. As shown in Fig. 6, our method still has some limitations. Since we derive alpha blending across shaders based on the Lambertian surface assumption, visible seams may occur at the boundaries between blocks on non-Lambertian surfaces, such as water surfaces. Combining physically-based rendering with multiple shaders blending may alleviate this problem. The deferred MLP in City-on-Web has limited representation ability for view-dependent color, which might cause numerous near-camera floaters. Taking the similar strategy used in [42, 12] by utilizing priors to pre-trim the scene or preprocessing the data are possible solutions.
Supplementary Material
7 Proof of 3D consistency
For a given sampling point $i$, suppose it is located within the region $\mathcal{K}$ of block $k$, which contains a total of $N_{k}$ sampling points. The diffuse color, feature, and opacity output by the shader of block $k$ are denoted as $\boldsymbol{c}^{k}_{i}$, $\boldsymbol{F}^{k}_{i}$, and $\alpha^{k}_{i}$ respectively. Then, for this ray, the output diffuse color $\boldsymbol{c}^{k}_{d}$ and specular feature $\boldsymbol{F}^{k}$of block $k$ are caculated as follows. $\boldsymbol{h}^{k}$ represents either the diffuse color or specular feature.
$\boldsymbol{h}^{k}(\boldsymbol{r})$ $=\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{h}_{i}^{k}$ (8)
$\alpha^{k}$ $=\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}$
By integrating the rendering results of each block’s shader through the volume rendering between blocks, the final diffuse color $\boldsymbol{c}_{d}$ and specular feature $\boldsymbol{F}$ of the ray can be obtained.
$\boldsymbol{h}(\boldsymbol{r})=\sum_{k}\prod_{j=1}^{k-1}(1-\alpha^{j})\cdot\boldsymbol{h}^{k}$ (9)
Note that
$\quad\ 1-\alpha^{k}$ (10)
$=1-\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}$
$=1-\alpha_{1}^{k}-(1-\alpha_{1}^{k})\alpha_{2}^{k}-(1-\alpha_{1}^{k})(1-\alpha_{2}^{k})\alpha_{3}^{k}-\cdots$
$=(1-\alpha_{1}^{k})(1-\alpha_{2}^{k}-(1-\alpha_{2}^{k})\alpha_{3}^{k}-\cdots)$
$=(1-\alpha_{1}^{k})(1-\alpha_{2}^{k})(1-\alpha_{3}^{k}-\cdots)$
[…]
$=\prod_{i=1}^{N_{k}}(1-\alpha_{i}^{k})$
Assuming the block positions are already depth-sorted, meaning there are $M$ sampling points along the ray, with each block $k$ containing $N_{k}$ sampling points, let $\alpha_{i}$ denote the $i$-th sampling point on the ray, and $\alpha_{i}^{k}$ denote the $i$-th sampling point in block $k$ along the ray, then
$\boldsymbol{h}(\boldsymbol{r})$ $=\sum_{k}\prod_{j=1}^{k-1}\prod^{N_{j}}_{i=1}(1-\alpha_{i}^{j})\cdot\boldsymbol{h}^{k}$ (11)
$=\sum_{k}\prod^{N_{1}+\cdots+N_{k-1}}_{i=1}(1-\alpha_{i}^{j})\cdot\left(\sum_{i=1}^{N_{k}}\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{h}_{i}^{k}\right)$
$=\sum_{k}\sum_{i=1}^{N_{k}}\prod^{N_{1}+\cdots+N_{k-1}}_{i=1}(1-\alpha_{i}^{j})\prod_{j=1}^{i-1}(1-\alpha_{j}^{k})\cdot\alpha_{i}^{k}\boldsymbol{h}_{i}^{k}$
$=\sum_{i=1}^{M}\prod^{i-1}_{i=1}(1-\alpha_{j})\alpha_{i}\boldsymbol{h}_{i}$
Eq. 11 shows that the diffuse color and specular feature we finally obtain by depth-sorting blocks and alpha blending along blocks are consistent with the results of volume rendering integration along the entire ray. Therefore, our method ensures the three-dimensional consistency of rendering.
Figure 7: Visulization of Campus Dataset.
8 Dataset
To demonstrate the effectiveness of our method, experiments are conducted on a variety of large scenes. The main experiments reported in this paper involve three types of environmental scene datasets of various scales. The Campus dataset is captured at a altitude of about 180 meters, covering an area of approximately 960,000 $m^{2}$. The Matrix City Dataset, captured at a altitude of 200 meters, is sparser than the Campus dataset, thus covering a larger area. The Mill 19 dataset covers a total of about 200,000 $m^{2}$. We adopt a circular data capture method for photographing, as shown in Fig. 7. We find that this method often results in a higher overlap rate, allowing for a more accurate estimation of camera poses. Our dataset was captured over 8 hours on a cloudy day, with a fixed exposure setting to ensure almost identical appearance of photos taken at different times. We used Colmap to estimate camera poses. Feature matching was done using a vocabulary tree, followed by a hierarchical mapper followed by a few iterations of triangulation and bundle adjustment to estimate camera poses.
Figure 8: Visulization of ou efficient non-overlapping partition strategy. The red square represents the $[-1,1]^{2}$.
9 Efficiency of Non-overlapping Partition Strategy
In the stage of segmenting the scene into distinct blocks, we initially rotate the scene to align it parallel with the xy-plane, then proceed to segment the entire space based on the xy coordinates of spatial points. This strategy’s merit lies in ensuring each segmented block is a bounded area. This is in contrast to methods like segmentation strategy of block-nerf  [34] and  [37], which cannot assure boundedness in the reconstructed area, potentially leading to substantial memory resource wastage as illustrated in the  Fig. 8. Our approach allows us to represent the same size area with a bounded region of $[-1,1]^{3}$, maintaining the same representation resolution, instead of using an unbounded region that contract to $[-2,2]^{3}$ like in the MERF  [30] and Mip-NeRF 360 [1]. Consequently, this enables the reduction of the resolution of the xy-plane from $4096^{2}$ to $2048^{2}$ without loss in performance. Thus we can effectively reduce the usage of VRAM, especially across the three planes.
10 Impletation Details.
For blocks at the boundaries of the entire scene, an unbounded scene representation is required to represent areas outside the block boundaries. We follow the same approach as MERF to compute ray-AABB intersections trivially. To be specific, we employ the scene contraction function to project the scene external to the unit sphere into a cube, which has a radius of 2. The definition of the $j-th$ coordinate for a contracted point is as follows:
$\operatorname{contract}(\mathbf{x})_{j}=\left\{\begin{array}[]{ll}x_{j}&\text{if }\|\mathbf{x}\|_{\infty}\leq 1\\ \frac{x_{j}}{\|\mathbf{x}\|_{\infty}}&\text{ if }x_{j}\neq\|\mathbf{x}\|_{\infty}>1\\ \left(2-\frac{1}{\left|x_{j}\right|}\right)\frac{x_{j}}{\left|x_{j}\right|}&\text{ if }x_{j}=\|\mathbf{x}\|_{\infty}>1\end{array},\right.$ (12)
11 More Results
We provide additional qualitative results and on Matrix City, Mill 19, Campus dataset as shown in Fig. 9.
12 Discussion
Recently, some research has also enabled real-time rendering of large scenes. UE4-NERF [13], building on the MobileNerf framework, divides large scenes into smaller segments for reconstruction and then renders the large-scale scene using the mesh rasterization pipeline. Like MobileNerf, UE4-NERF begins with a $128^{3}$ grid, which assumes an even distribution of scene details in all directions. However, data captured through oblique photography often appears ’flat’, meaning there’s dense information when projected onto the xy plane, but sparser in the vertical direction due to mostly empty areas. Therefore, UE4-NERF needs more segments for large-scale reconstruction to ensure an even distribution of details in all directions in a block. For example, their Construction Site scene required about 40 blocks to reconstruct a $420m^{2}\times 240m^{2}$ area, leading to significant memory and VRAM usage, approximately 25GB. Even with UE4’s dynamic mesh-based loading, VRAM usage is around 11GB. Such high payload and VRAM demands make it challenging to extend this technology to web platforms and consumer-grade GPUs.
Additionally, NeuRas [3] has also made progress in real-time rendering of large scenes. It uses texture-less geometry from already reconstructed large scenes. NeuRas applys feature map textures and combines mesh rasterization results with view direction to query a small MLP for view-dependent colors. They optimize the feature texture and view-dependent MLP to enhance rendering results. However, this method also suffers from high memory usage. Our experiments show that obtaining texture-less geometry with ContextCapture [^1] https://www.bentley.com/software/itwin-capture-modeler/ for the Campus scene requires about 6GB, and surface reconstruction via neural rendering methods needs approximately 10GB without mesh simplification.
Both these methods have achieved excellent results in real-time rendering of large scenes, but their significant memory and VRAM requirements limit their expansion to web platforms. Our method can implement real-time rendering of large scenes with a payload and VRAM consumption acceptable for web platforms and consumer-grade graphics cards.
Figure 9: Qualitative visualization of geometry and rendering result on five datasets.